OpenStudy (anonymous):

How do you find the curl of function f(x,y)=1111+(x/1100)^2

OpenStudy (anonymous):

\[curl f=\left[\begin{matrix}i & j & k \\ \frac{\partial }{\partial x}& \frac{\partial }{\partial y}& \frac{\partial }{\partial z}\\ f_{x} & f_{y}& f_{z}\end{matrix}\right]\]

OpenStudy (anonymous):

in your case f looks like scalar field. So there is no curl

OpenStudy (anonymous):

is that fx , fy and fz in 3rd row are you sure? isn't it cifficient of i j and k?

OpenStudy (anonymous):

that's what fx , fy and fz are. They are the components of f

OpenStudy (anonymous):

f should be a vector, if you whant to find it's curl. So it so it should look somthing like this: \[f(x,y)=f_{x} i +f_{y} j\]

OpenStudy (anonymous):

Here f describes pressure vector field in 2 dimension

OpenStudy (anonymous):

When you make contour plot you can see, pressure increasing from higher to lower region

OpenStudy (anonymous):

DO you know how to write this equation as vector field? I am given only this equation

OpenStudy (anonymous):

as I said befor. This equation asigns a numerical (scalar) value to every point in xy plane. So there is no vector field. The only vector field asociated with this sclara field is the field of grad f. But the curl(grad f)=0 always. That's all there is about this.

OpenStudy (anonymous):

ok thnks